OpenStudy (anonymous):

\[\LARGE \color{green}{\star}\text{Happy 13th birthday Parth} \color{red}{\star}\] We hope you have a great day and celebrate well!!

OpenStudy (anonymous):

That didn't go as planned... It was supposed to be pretty Latex :D xD Apoorvk sends his biggest wishes and hopes you have a really, really fun time! Welcome to the Teenage years! :D

OpenStudy (anonymous):

\[\ \text \ happy \ 13th \ birthday\ Parth! \]

OpenStudy (zepp):

\[\LARGE \color{green}{\star}\text{Happy 13th birthday Parth} \color{red}{\star}\]

OpenStudy (anonymous):

Show off ^

OpenStudy (zepp):

^_^

OpenStudy (anonymous):

Happy 556th PARTH! :D

OpenStudy (anonymous):

Happy birthday Parth from all the admins at OpenStudy :)

OpenStudy (anonymous):

Who changed meh post?

OpenStudy (turingtest):

me

OpenStudy (turingtest):

want it back the old way?

OpenStudy (anonymous):

Guilty party haha is there a way to make it such that Zepp can have credit for creating it then? He made that. lol

OpenStudy (zepp):

Who copied my \(\LaTeX\)? :P

OpenStudy (anonymous):

Turing did it XD

myininaya (myininaya):

Happy Bday!

OpenStudy (anonymous):

Thanks :) All credit for the birthday sign go to @Zepp :) lol zepp do you wanna delete our unnecessary posts?

OpenStudy (anonymous):

happy birthday ,you're a man now

hero (hero):

HBD

OpenStudy (diyadiya):

\[\Large \mathcal {HaPpY~BiRtHdAy~PaRtH } =)\]

OpenStudy (lgbasallote):

hah finally you're not a 12 year old "genius" >:)))) lol just kidding. happy birthday from all the lgbaists

OpenStudy (karatechopper):

HAAAAPPPPPYYYYY BBBIIIIRRRTTHHDDDAAAYYYYYYY!!!!! Sorry i dont know how to do the latex:D

Parth (parthkohli):

Woah! The best birthday ever? Fo' sho'!

OpenStudy (anonymous):

BECAUSE OF ME! :D

Parth (parthkohli):

And I don't care as far as the LaTeX is concerned.

Parth (parthkohli):

Yeah, @Pac1f1cIslander :)

OpenStudy (anonymous):

:D All grown up haha :D And seriously don't forget Apoorvk. He really wanted it to be special for you and sent his best wishes and hopes you have an incredible day. And yes, I apologize about my lack of latex abilities.

OpenStudy (anonymous):

I knew it! xD

Parth (parthkohli):

Yeah.. I talked to apoorv last night. :)

OpenStudy (karatechopper):

Yes. Apoorvk had sent mee all his regards for u! So.... HAAPPPYYY BBIIRRTTHHDDAYY!! -Apoorvk

Parth (parthkohli):

Woh toh sahi hai, lekin KC aapko school nahi jaana?

OpenStudy (lgbasallote):

\[\textbf{marvel at the } \mathbf{\LaTeX}\]@rebeccaskell94

OpenStudy (karatechopper):

Parth meri abhi chootiya chal rahi hai!

OpenStudy (anonymous):

I always do LG xD

Parth (parthkohli):

\( \Huge \color{Maroon}{\mathtt{\text{I can live without} \LaTeX}} \)

OpenStudy (karatechopper):

lol, BTW. I am wonderin when apoorvk will be back.:( miss him already

Parth (parthkohli):

Of course :/

OpenStudy (karatechopper):

He have an email he ever reply to?

OpenStudy (lgbasallote):

\[\Huge \color{midnightblue}{\ddot \smile}\]

OpenStudy (anonymous):

I have an idea!! Brb

OpenStudy (karatechopper):

I have an imposter in chat?

Parth (parthkohli):

You still won't be able to implement it in the next 9 hours, because I am going to school in a few.

OpenStudy (lgbasallote):

\[\Huge \stackrel{\LaTeX}{\stackrel{\LaTeX}{\stackrel{\LaTeX}{\LaTeX}}}\]

OpenStudy (anonymous):

psh show off

Parth (parthkohli):

LaTeX LaTeX LaTeX LaTeX Why LaTeXify things that you can type out right on the keyboard?

OpenStudy (karatechopper):

what code?

OpenStudy (lgbasallote):

\[\Huge \stackrel{\text{mada}}{\stackrel{\mathsf{mada}}{\mathbb{dane}}}\]

OpenStudy (lgbasallote):

\[\huge \begin{array}{cccc} \mathsf{you} \\ \mathcal{don't} \\ \mathbf{think} \\ \mathbb{fourth-dimensionally} \end{array}\]

OpenStudy (anonymous):

Darn. I was trying to find an old Latex Apoorvk did for someone's birthday and then I was going to put it over here so it would be lovely. But I couldn't find it. Sorry, Parth. Happy Birthday, either way. :) Hope you have fun joining the rest of us in teenage-hood :D

Parth (parthkohli):

\[\begin{array}{} \mathtt{I} \\ \mathbf{Can} \\ \mathbf{Do} \\ That \\ Too! \end{array} \]

OpenStudy (lgbasallote):

\[\huge \begin{array}{cccc} \color{blue}{\mathbf{don't}} \\ \color{red}{\mathtt{do}} \\\color{gold}{\mathbb{this}} \\ \color{green}{\mathsf{Kohli}} \end {array}\]

OpenStudy (lgbasallote):

dont challenge me...it'll only end badly

OpenStudy (anonymous):

\\[\text this\ is\ all \ I \ can\ do \]

Parth (parthkohli):

\[ \begin{array}{} \text{Too busy to start a }\LaTeX \textbf{ fight}. \end{array} \]

OpenStudy (lgbasallote):

\[\LARGE \mathtt{\color{green}{g} \color{red}{o} \color{darkblue}{o} \color{orange}{d}}\]

OpenStudy (lgbasallote):

sheesh..im out of touch with my artistic side...these color combos suck

OpenStudy (anonymous):

Agreed.

OpenStudy (anonymous):

Happy birthday party:D

OpenStudy (anonymous):

\[\huge≤\color{red}{\textbf{Happy}} \space \color{blue}{Birthday} \space \color{green}{\text{Parth!}}≥\]Sorry, I don't know as much \(\LaTeX\) as some people XD

OpenStudy (anonymous):

My gift to Parth :D http://cdn.memegenerator.net/instances/400x/17836797.jpg

OpenStudy (anonymous):

@ParthKohli @karatechopper aap desi hai?

OpenStudy (ash2326):

\[\Huge\color{gold} \bigstar\color{Green} { \mathbb {Happy Birthday Parth :)}}\]

OpenStudy (waheguru):

≤Wow, Parth is just 13 and has done so much, happy birthday bro!≥

OpenStudy (waheguru):

Now How Do I make that in latex

OpenStudy (karatechopper):

@sarah43 YES!

mathslover (mathslover):

Happy birthday parth So many wishes !!! may god bless u my friend !!

OpenStudy (anonymous):

It may not be your bday there, but its still your bday in america, so HAPPY BDAYYY!!!!! hahaha. Hope all your wishes come true!

Parth (parthkohli):

:)

OpenStudy (anonymous):

\[\huge\mathfrak {\color{#DD721F}{Happy}} \text{ } \mathfrak{\color{#1FA1DD}{birthday!!!}} \text{ } \mathsf{\color{green}{@ParthKohli}}\] \[\large \text{**}celebrating\text{ } 13 \text{ } happy \text{ } years \text{ }of \text{ }awesomeness \text{ }\text{**}\]

Parth (parthkohli):

Apoorv?